5) A mortgage company classifies its borrowers into three categories: Low Risk, Medium Risk, and High Risk. From experience, the company knows that:

  • 3% of low risk borrowers eventually default on their mortgages.
  • 9% of medium risk borrowers eventually default on their mortgages.
  • 13% of high risk borrowers eventually default on their mortgages.

The mortgages for 129 high risk borrowers are put together into one portfolio. The company determines that they will profit on the portfolio as long as no more than 20% of borrowers with mortgages in the portfolio default. What is the probability that the company makes a profit on the portfolio?

Round your answer to 4 decimal places.
Answer #1
for normal distribution z score =(p̂-p)/σp
here population proportion=     p= 0.1300
sample size       =n= 129
std error of proportion=σp=√(p*(1-p)/n)= 0.0296

P( no more than 20% of borrowers with mortgages in the portfolio default ):

probability =P(X<0.2)=(Z<(0.2-0.13)/0.03)=P(Z<2.36)=0.9909